# Integration of TinyML and LargeML: A Survey of 6G and Beyond

### v1

@@ III-C6 Efficient ISAC
ISAC is a key 6G requirement, unifying advanced sensing, communication, and precise positioning within a unified framework [85]. ISAC supports critical functions such as range, velocity, and angle estimation, as well as object detection, localization, imaging, and mapping. This integration reduces redundancy, optimizes resource use, and boosts overall network performance. High-precision sensing is essential for latency-sensitive applications like autonomous vehicles, smart cities, and AR, all of which demand real-time, location-specific data. Positioning accuracy in 6G is targeted within $1-10$ cm [81]. The integration of TinyML and LargeML can significantly enhance the efficiency, accuracy, and robustness of ISAC systems. For example, TinyML enables on-device processing and feature extraction, allowing initial estimation of key parameters and basic object detection. This local filtering reduces noise and improves signal fidelity. Processed outputs are then sent to LargeML systems for deeper analysis and validation, improving detection accuracy and reducing false positives. Additionally, TinyML supports real-time mapping by continuously updating environmental models, which is crucial for autonomous navigation that relies on high-resolution, up-to-date spatial information.
@@ IV-A Transfer Learning
TL, a subset of ML methods, focuses on reusing the knowledge from pre-trained models across related tasks and data domains. The technique provides an effective approach to TinyML–LargeML integration in 6G and beyond networks by combining the strengths of server-based large-scale training and on-device adaptability [86]. Fig. 4 depicts a TL-based two-way integration framework, consisting of three main steps.
@@
Server-side training: A large dataset of $D_{S}$ samples, denoted ${{\cal D}_{S}}=\left\{{1,\ldots,{D_{S}}}\right\}$, is used to train a LargeML model, denoted $f_{S}$, on a high-performance cloud or edge server, referred to as the source domain. Let ${\bf{\Theta}}=\left\{{{{\bf{w}}_{l}},{{\bf{b}}_{l}}}\right\}_{l=1}^{L}$ be the set of parameters comprising all weights ${{{\bf{w}}_{l}}}$ and biases ${{{\bf{b}}_{l}}}$ across $L$ layers, which form the model pipeline. The training objective is to minimize the loss function, i.e.,
${\bf{\Theta}}=\arg\mathop{\min}\limits_{\bf{\Theta}}\frac{1}{{{D_{S}}}}\sum\nolimits_{i\in{{\cal D}_{S}}}{{\cal L}\left({{{\bf{y}}_{i}},{f_{S}}\left({{{\bf{x}}_{i}}}\right)}\right)},$ (1)
where ${\left\{{{{\bf{x}}_{i}},{{\bf{y}}_{i}}}\right\}_{i\in{{\cal D}_{S}}}}$ represents the training samples, and ${{f_{S}}\left({{{\bf{x}}_{i}}}\right)}$ denotes the model’s activation after processing input feature ${\bf{x}}_{i}$. The loss function ${\cal L}(\cdot)$ varies depending on the learning task and may include metrics such as cross-entropy, mean squared error, MAE, or log-likelihood. For example, the loss function of a linear regression problem over dataset $\left({{\bf{X}},{\bf{y}}}\right)$ is given by ${{\cal L}_{0}}\left({{\bf{w}}|{\bf{X}},{\bf{y}}}\right)=\frac{1}{M}\sum\nolimits_{m=1}^{M}{{{\left({{{\bf{w}}^{T}}{{\bf{x}}_{m}}-{y_{m}}}\right)}^{2}}}$, where ${{{\bf{x}}_{m}}}$ and $y_{m}$ are the input feature vector and label scalar, respectively, of the data sample $m$ among $M$ data points of $\left({{\bf{X}},{\bf{y}}}\right)$. The server’s high computational resources and storage capabilities allow it to process large datasets and train complex models effectively.
Server-side training can adopt either a standalone architecture (e.g., two-tiered client-edge, two-tiered client-cloud) or a multi-level collaboration model [87]. Compared to standalone setups, multi-level collaboration offers greater flexibility in balancing communication and computational overheads. Four representative scenarios include: (i) Hybrid training, where initial training of LargeML is conducted in the cloud, followed by deployment to edge servers for fine-tuning and real-time adaptation using local data; (ii) Distributed fashion, where the training workload is distributed between the cloud and edge servers, with large-scale, resource-intensive tasks handled in the cloud and latency-sensitive, less intensive tasks processed at the edge; (iii) Hierarchical processing, where edge servers perform data preprocessing to reduce data volume and complexity before transmitting it to the cloud for deeper analysis, while updates from the cloud can subsequently relayed back to edge servers for local execution and adaptation; and (iv) Enhanced availability, wherein hierarchical server models improve system resilience by allowing edge servers to operate independently during cloud outages.
@@
Pre-trained knowledge transfer: After training on the source domain, selected components of the LargeML $f_{S}$, specifically a subset of ${\bf{\Theta}}$, are transferred to the target domain (i.e., TinyML devices). These transferred components encapsulate the general features and patterns learned from ${\mathcal{D}}_{S}$. Effective knowledge transfer in TL hinges on four primary considerations as described below [88, 89, 90].
What to transfer
The “what” aspect refers to identifying the source model elements to be reused in the target domain. These may include the components or knowledge that are likely to benefit the target task. We denote $t_{S}$ as the transferred model, where $t_{S}$ is a subset of ${\bf{\Theta}}$. This decision is the most critical in TL.
When to transfer
The “when” aspect involves determining the appropriate timing and frequency of knowledge transfer between the source and target domains. It requires the selection of a strategy for updating the target domain with knowledge from the source domain. Strategies can be grouped into three types [91]: (i) Real-time TL: Knowledge transfers occur immediately after updates in the source domain, which is suitable for dynamic environments requiring rapid adaptation; (ii) Non-real-time or semi-real-time TL: Knowledge transfers follow a fixed schedule (e.g., daily, weekly) or occur at defined operational checkpoints, which suits stable environments or models where immediate updates are not critical; and (iii) On-demand TL: Knowledge transfers occur only when explicitly requested by the target domain, which benefits scenarios in which the target domain operates autonomously and requires updates under specific conditions only.
Where and how to transfer
@@
The “where” aspect refers to identifying the target domain where the knowledge will be applied (e.g., TinyML devices). The “how” aspect ensures compatibility and maximizes the effectiveness of knowledge use in the target domain. Effective transfer depends on both domain similarity and the selection of optimization techniques. Transfer is typically more successful between domains with overlapping characteristics. For example, a resource management model trained on 5G network infrastructure may be adapted to monitor and optimize components in a 6G network [92]. The hardware and software constraints of TinyML devices also affect both where and how knowledge transfer occurs. The transferred model must be compatible with these constraints to function efficiently under limited computational and memory resources.
Examined through these guiding questions, TL can be categorized according to several criteria [89]. One approach focuses on label availability: (i) In transductive TL, labeled data is available only in the source domain, and the source and target tasks are similar; (ii) In inductive TL, labeled data is available only in the target domain, and the source and target tasks differ; and (iii) In Unsupervised TL, neither domain has labeled data, and the target task is distinct yet related to the source task. Another categorization approach depends on the alignment of feature and label spaces between the source and target domains. When both domains share identical features and label spaces, the setting is classified as homogeneous TL; otherwise, it is considered non-homogeneous TL. A third, widely used classification is based on the transfer mechanism. Specifically, Instance-based TL selects and re-weights specific instances (data points) from the source domain to improve their relevance in the target domain, aiming to identify the most useful source data points for training in the target domain [93]. Feature-based TL seeks to identify shared feature representations for the source and target domains [94]. By transforming features from both domains into a shared feature space, using techniques such as dimensionality reduction, feature extraction, or transformations, this approach facilitates more effective knowledge transfer. Parameter-based TL reuses model parameters (e.g., weights and biases) learned from the source domain and transfers them to the target domain [95]. Typically, a model is trained on a large source-domain dataset, and the learned parameters serve as an initialization point for training on the target dataset, enabling the model to retain useful knowledge while adapting to the new domain. Finally, Relational-based TL is used when data is structured as relationships or graphs (i.e., learned rules) [96]. It transfers structural or relational information, such as node embeddings, relational patterns, or graph kernels, from the source domain to support learning in the target domain.
@@
Model personalization and on-device adaptation: Consider a set of $N$ TinyML devices in the target domain, denoted ${\cal N}=\left\{{1,\ldots,N}\right\}$. Each device ${n\in{\cal N}}$, equipped with limited computational resources and storage, adapts a pre-trained model $t_{S}$ along with its TinyML model $g_{n}$, using its individual, smaller, scenario-specific dataset ${{\cal D}_{n}}=\left\{{1,\ldots,{D_{n}}}\right\}$. This adaptation involves fine-tuning (i.e., freezing $t_{S}$ while updating the parameters associated with $g_{n}$) and additional training to tailor the model for relevance to the device’s specific tasks. This process is referred to as model personalization. The concept of an adaptation layer introduced in [97] augments this process by minimizing the general loss function and the maximum mean discrepancy (MMD) loss. MMD measures the divergence between source and target domain distributions by evaluating their distance in a reproducing kernel Hilbert space (RKHS). It serves as a criterion for comparing distributional discrepancy in transfer learning. An empirical estimate of the MMD loss function between ${\cal D}_{S}$ and ${\cal D}_{n}$ is given by ${{\cal M}_{{{\cal D}_{S}},{{\cal D}_{n}}}}=\big{\|}\frac{1}{{{D_{S}}}}\sum\nolimits_{i\in{{\cal D}_{S}}}{\phi\left({{{\bf{x}}_{i}},{{\bf{y}}_{i}}}\right)}-\frac{1}{{{D_{n}}}}\sum\nolimits_{j\in{{\cal D}_{n}}}{\phi\left({{\bf{x}}_{j}^{n},{\bf{y}}_{j}^{n}}\right)}\big{\|}_{\cal H}$ [97], where $\phi\left(\cdot\right)$ is the nonlinear mapping into RKHS $\cal H$, and ${\left\{{{\bf{x}}_{j}^{n},{\bf{y}}_{j}^{n}}\right\}_{j\in{{\cal D}_{n}}}}$ represents the samples from ${\cal D}_{n}$. Consequently, the objective of minimizing the loss function for local training on client ${\left\{n\right\}_{n\in{\cal N}}}$ is expressed as
${{\bf{\Theta}}_{n}}=\arg\mathop{\min}\limits_{{{\bf{\Theta}}_{n}}}\frac{1}{{{D_{n}}}}\sum\limits_{j\in{{\cal D}_{n}}}{{\cal L}\left({{\bf{y}}_{j}^{n},{t_{S}}\left({{\bf{x}}_{j}^{n}}\right)}\right)}+\lambda{\cal M}_{{{\cal D}_{S}},{{\cal D}_{n}}}^{2},$ (2)
where $\lambda>0$ is the trade-off parameter. However, given the simplicity of this shallow model, its performance is limited. To further mitigate domain shift between the source and target domains, correlation alignment (CORAL) can be applied [98]. The CORAL loss between ${\cal D}_{S}$ and ${\cal D}_{n}$ is formally defined as ${\ell_{{{\cal D}_{S}},{{\cal D}_{n}}}}=\frac{1}{{4{d^{2}}}}\left\|{{{\bf{R}}_{{{\cal D}_{S}}}}-{{\bf{R}}_{{{\cal D}_{n}}}}}\right\|_{F}^{2}$, where $d$ is the dimension of embedding features, ${\left\|\cdot\right\|_{F}}$ denotes the Frobenius norm, and ${{\bf{R}}_{\cal D}}$ denotes the covariance matrix of dataset $\cal D$. Consequently, (2) can be reformulated as
${{\bf{\Theta}}_{n}}=\arg\mathop{\min}\limits_{{{\bf{\Theta}}_{n}}}\frac{1}{{{D_{n}}}}\sum\limits_{j\in{{\cal D}_{n}}}{{\cal L}\left({{\bf{y}}_{j}^{n},{t_{S}}\left({{\bf{x}}_{j}^{n}}\right)}\right)}+\lambda{\ell_{{{\cal D}_{S}},{{\cal D}_{n}}}}.$ (3)
Since TinyML devices can perform tasks locally after knowledge transfer,they retain sensitive data on-device, eliminating the need for continuous cloud transmission or inter-device sharing, thus improving privacy.
Recent studies have explored the integration of TinyML and LargeML through TL for various applications [38, 36, 99, 100, 101, 102, 103]. For example, [38] proposes a TL-based system which combines TinyML and LargeML, named TinyTL. The framework employs ProxylessNAS-Mobile as the backbone NN for source-domain pre-training, and fine-tunes the model on TinyML devices using a pre-trained once-for-all network. The parameter-based TL approach transfers feature extractor weights. Target tasks include object and facial attribute classifications, such as recognition of cars, aircraft, flowers, birds, pets, food, and celebrities. [36] introduces a bidirectional integration framework incorporating parameter-based TL. The system pre-trains a CNN in the source domain, then adapts the CNN to TinyML devices using Tensorflow Lite Micro with pruning and quantization techniques. Target tasks include presence detection and classification inside cars. In [99], the authors developed a bidirectional integration system using feature-based TL aimed at object classification and human activity recognition for IoT health tracker applications. The source domain backbone is a deep neural network (DNN), and the target model runs on CMSIS-NN.
@@
Building on prior work, the study in [100] proposes a feature-based TL system that exploits VGG-16, DenseNet, and MobileNet as backbone architectures for source-domain pre-training. For deployment in the TinyML domain, the system uses the EON Tuner tool, integrated with the Edge Impulse platform, to perform face mask detection and compare performance across the three models. Another study [101] applies a parameter-based TL approach for soil humidity prediction in smart agriculture. The model combines a DNN and long short-term memory (LSTM) architecture for source-domain pre-training. In the target domain, fine-tuning with TensorFlow Lite Micro and quantization techniques are used for deployment on TinyML devices. In [102], the authors introduce a feature-based TL bidirectional integration system for motion classification. The approach uses principal component analysis for dimensionality reduction and trains a CNN-LSTM model on large-scale, diverse source-domain datasets. Fine-tuning is performed on TinyML devices, with a Raspberry Pi as an edge intermediary platform. A feature-based TL system presented in [103] incorporates meta-learning for a TinyML–LargeML framework and cross-domain classification tasks. The system uses a DNN pre-trained with MCUNet, MobileNetV2, and ProxylessNAS architectures in the source domain and performs fine-tuning on TinyML devices using a task-adaptive sparse-update method. The design prioritizes adaptability to dynamic user contexts and device constraints, using few-shot learning with minimal samples.
IV-B Federated Transfer Learning
@@ IV-B1 Vanilla Federated Transfer Learning
The TL strategies described in Section IV-A are effective when source and target domains share related tasks and datasets. However, their applicability in 6G environments is limited due to the prevalence of highly non-IID data across massively connected devices. FTL, which combines the strengths of FL and TL [104], is better suited for such heterogeneous settings. In 6G’s device-dense landscape, effective personalization is crucial for optimizing performance on device-specific tasks. Notably, FL enables collaborative training across distributed TinyML systems while preserving data locality on resource-constrained devices [105]. FTL enhances this by distinguishing between source clients (who participate in training the initial global model) and target clients (who personalize the model to their specific needs). As both FL and TL strategies emphasize privacy, FTL effectively addresses the dual challenges of data heterogeneity and privacy in integrated TinyML–LargeML systems. Fig. 5 illustrates the workflow of this FTL-based approach, as outlined below.
The FTL process begins with the standard FL protocol, differing primarily in the initialization of the global model. Specifically, the FL phase includes the following steps.
@@
    Initial global model sharing: In contrast to conventional FL, where the server initializes the global model with random weights, FTL initializes the global model using a pre-trained LargeML model $f_{S}$, trained on the server’s large dataset ${\cal D}_{S}$. As with (1), the objective function for training is given by $f_{S}^{\left(0\right)}=\arg\mathop{\min}\nolimits_{{f_{S}}}\frac{1}{{{D_{S}}}}\sum\nolimits_{i\in{{\cal D}_{S}}}{{\cal L}\left({{{\bf{y}}_{i}},{f_{S}}\left({{{\bf{x}}_{i}}}\right)}\right)}$. Once trained, $f_{S}^{\left(0\right)}$ is distributed to a source client set ${\cal K}=\left\{{1,\ldots,K}\right\}$.
@@
    Local training: In the training round $t$, each source client ${k\in{\cal K}}$ uses its local dataset ${{\cal D}_{k}}=\left\{{1,\ldots,{D_{k}}}\right\}$ to train and update a local copy of the model. Local training enables the model to learn patterns and insights specific to each client’s data. Mathematically, the objective for minimizing the loss function during client $k$’s local training is expressed as $f_{S,k}^{\left(t\right)}=\arg\mathop{\min}\nolimits_{f_{S}^{\left({t-1}\right)}}\frac{1}{{{D_{k}}}}\sum\nolimits_{j\in{{\cal D}_{k}}}{{\cal L}\big{(}{{\bf{y}}_{j}^{k},f_{S}^{\left({t-1}\right)}\left({{\bf{x}}_{j}^{k}}\right)}\big{)}}$.
@@
    Local model update: After local training, each source client uploads its updated model parameters $f_{S}^{\left(t\right)}$ to the server for aggregation.
@@
    Server-side aggregation: The server obtains global model parameters by minimizing the global loss function $f_{S}^{\left(t\right)}=\arg\mathop{\min}\nolimits_{f_{S,k}^{\left(t\right)}}\sum\nolimits_{k\in{\cal K}}{\frac{{{r_{k}}}}{{{D_{k}}}}\sum\nolimits_{j\in{{\cal D}_{k}}}{{\cal L}\big{(}{{\bf{y}}_{j}^{k},f_{S,k}^{\left(t\right)}\left({{\bf{x}}_{j}^{k}}\right)}\big{)}}}$ where ${r_{k}}={D_{k}}/\sum\nolimits_{k=1}^{K}{{D_{k}}}$ represents the ratio of client $k$’s data size to the total data size. This aggregation problem can be solved using the FedAvg algorithm [106], a proportion weighting strategy [107], or more advanced methods that account for data heterogeneity and varying client contributions.
@@
Step 6 presents a knowledge transfer phase. Once the global model has been sufficiently trained and demonstrates strong generalization, the server transfers the useful components of the pre-trained global model $f_{S}$, denoted $t_{S}$, to the target TinyML devices (target clients). This transferred model, derived form the pre-trained global model, encapsulates the knowledge learned from source clients and serves as the foundation for the next phase.
Step 7 stands for model personalization and on-device adaptation. Target clients typically hold smaller datasets or domain-specific datasets that differ from those of source clients. To tailor the transferred model $t_{S}$ to local tasks, each target client performs personalization through TL. We consider the set of target TinyML devices ${\cal N}=\left\{{1,\ldots,N}\right\}$. Each target client fine-tunes $t_{S}$ with its local dataset ${\cal D}_{n}$. Depending on the similarity of the global model’s training data and the target client’s data, fine-tuning may involve updating the entire model $t_{S}$ or only a subset of layers. To support deployment in resource-constrained environments, each target client employs a TinyML model $g_{n}$. The personalized model is then trained using either objective functions (2) or (3). The adaptation process harnesses the advantages of TL, enabling target clients to attain high performance without the need to train a model from scratch. Consequently, this approach conserves computational resources and reduces training time.
@@
Numerous recent technical publications have explored the use of FTL to enable seamless integration between TinyML and LargeML across diverse application domains [108, 109, 110, 111]. For example, the work in [108] introduces TinyFedTL, an open-sourced, parameter-based FTL framework designed for TinyML–LargeML systems. The framework pre-trains MobileNetV2 on a source domain and applies Perf-MobileNet, part of the TinyML Perf Benchmark, for personalized model adaptation on IoT edge devices. Experimental results validate the effectiveness of TinyFedTL in enhancing the quality of experience (QoE) under resource constraints, without compromising data privacy. Proposed in [109], FedHealth is a parameter-based FTL framework for wearable healthcare devices. FedHealth employs a CNN backbone network for source-domain pre-training, followed by fine-tuning during the TL phase to adapt the model to resource-constrained IoT devices. Evaluations of wearable activity recognition and auxiliary Parkinson’s disease diagnosis tasks demonstrate that FedHealth achieves high accuracy, without compromising privacy and security. A parameter-based FTL framework incorporating auxiliary classifier generative adversarial networks, termed ACGAN-FTL, is proposed in [110]. ACGAN-FTL runs personalized models on Raspberry Pi-based TinyML devices and is designed to predict the quality of pre-baked carbon anodes in industrial production. Recently, the authors in [111] demonstrated the superior performance of FTL compared to standalone FL, TL, and Tensorflow Lite-based fusion approaches in both classification and regression tasks in TinyML–LargeML systems. The TL principle applied in this framework centers on feature-based TL, using Genann [112] as the backbone neural network for source-domain pre-training, followed by fine-tuning for device-specific adaptation across multiple TinyML platforms, including Arduino WiFi Rev2, Arduino MKR1010, ESP8266, and ESP32.
@@
From a hierarchical FTL perspective, the study in [113] introduces IoTDefender, an intrusion detection framework designed for resource-constrained IoT networks. IoTDefender employs three edge servers that run standard FL protocols on distinct IoT networks. Their local models are subsequently aggregated at a cloud server. IoTDefender uses a CNN backbone for source-domain pre-training, with subsequent fine-tuning for parameter-based TL and on-device adaptation in resource-constrained IoT devices. Expanding on the topology proposed in [113], the study in [114] proposes HFTL, a hierarchical FTL framework that incorporates additional fog servers. These fog servers aggregate models from associated edge servers before forwarding to a cloud server for further aggregation. HFTL uses VGG-16, ResNet50V2, and MobileNet as pre-trained models, followed by fine-tuning with an additional DNN for secure and efficient fault classification in additive manufacturing. The framework adopts parameter-based TL as its core approach.

### v2

@@ III-C6 Efficient ISAC
ISAC represents a defining requirement of 6G, merging communication, sensing, and localization into a unified framework [103]. TinyML supports on-device feature extraction and preliminary sensing tasks, while LargeML performs advanced analysis for object detection, imaging, and mapping. This integration enhances precision, reduces false positives, and enables real-time navigation and environmental awareness, being key for autonomous systems and smart cities.
III-D Discussions and Lessons Learned
The transition from 5G to 6G creates significant opportunity to integrate TinyML and LargeML to enhance AI capabilities in next-generation networks. Section III-A discusses the key challenges of TinyML, while Section III-B examines the limitations of LargeML. As highlighted in Section III-C, a hybrid framework that combines TinyML and LargeML leverages the strengths of distributed and centralized learning, forming a promising paradigm aligned with the 6G vision. In essence, TinyML enables edge-level data preprocessing to reduce bandwidth consumption and latency, whereas LargeML performs deeper analysis for high-accuracy insights. This bidirectional integration, powered by emerging technologies, helps fulfill the six core requirements of 6G networks. Detailed discussions of specific integration approaches are provided in Section IV.
@@ IV-A Transfer Learning
TL, a subset of ML methods, enables knowledge reuse from pre-trained models across related tasks and domains (readers can find more details of this technique in [178]), providing an effective mechanism for TinyML–LargeML integration in 6G networks. By combining large-scale server-based training with lightweight on-device adaptation, TL supports both scalability and personalization. A generic TL framework for integration consists of three main stages (see Fig. 4).
@@
Server-side training: A LargeML model $f_{S}$ is trained on a source-domain dataset ${\cal D}_{S}$} using high-performance cloud or edge servers. The objective is to optimize model parameters comprising weights and biases by minimizing the loss function
${f_{S}}=\arg\mathop{\min}\limits_{f_{S}}\frac{1}{{{|{\cal D}_{S}|}}}\sum\nolimits_{i\in{{\cal D}_{S}}}{{\cal L}\left({{{\bf{y}}_{i}},{f_{S}}\left({{{\bf{x}}_{i}}}\right)}\right)},$ (1)
[…]
Different server-side training architectures can be adopted [157], including: (i) hybrid training (cloud pre-training with edge fine-tuning), (ii) distributed training (workload split between cloud and edge), (iii) hierarchical processing (edge pre-processing before cloud training), and (iv) enhanced availability (separating edge servers during cloud outages). Compared to standalone setups (e.g., two-tiered client-edge/client-cloud), multi-level collaboration offers greater flexibility in balancing communication and computational overhead.
@@
Pre-trained knowledge transfer: After training in the source domain, selected model components $t_{S}\subset f_{S}$ are transferred to the target domain. The effectiveness of this process relies on answering four guiding questions [71, 127]:
  • What to transfer: The transferred knowledge may include model parameters (weights and biases), feature extractors, embeddings, or even relational structures. The choice depends on the similarity between source and target tasks. Parameters are most effective when tasks are closely related, while feature- or relational-based knowledge suits domains with shared patterns but different label spaces.
  • When to transfer: Knowledge can be transmitted within [128]: (i) Real-time, immediately after model updates, enabling rapid adaptation in dynamic environments; (ii) Periodic/semi-real-time, following scheduled intervals (e.g., daily, weekly), useful in stable or predictable contexts; and (iii) On-demand, initiated by the TinyML device when additional knowledge is required, reducing communication overhead in autonomous scenarios.
@@
  • Where and how to transfer: The “where” aspect refers to identifying the target domain where the knowledge will be applied (e.g., TinyML devices). The “how” aspect ensures compatibility and maximizes the effectiveness of knowledge use in the target domain. Effective transfer depends on both domain similarity and the selection of optimization techniques. Transfer is typically more successful between domains with overlapping characteristics.
Examined through these guiding questions, TL can be categorized according to several criteria [127]:
  • Label availability: Transductive (labels only in source), inductive (labels only in target), or unsupervised (no labels in either domain).
  • Domain alignment: Homogeneous (shared feature and label spaces) and non-homogeneous (otherwise).
  • Transfer mechanism: It includes approaches: (i) Instance-based TL, re-weighting or selecting relevant source samples; (ii) Feature-based TL, learning shared representations via dimensionality reduction or transformations; (iii) Parameter-based TL, reusing pre-trained weights and biases as initialization for target training; and (iv) Relational-based TL, transferring structural knowledge such as graphs, relationships, or relational patterns.
@@
Model personalization and on-device adaptation: Each TinyML device $n\in{\cal N}=\left\{{1,\ldots,N}\right\}$ adapts the transferred model using its local dataset ${\cal D}_{n}$}. Fine-tuning typically involves freezing pre-trained layers $t_{S}$ while updating TinyML lightweight layers $g_{n}$ for device-specific tasks. Domain adaptation methods, such as adaptation layers [51] or correlation alignment (CORAL) [156], help reduce distributional shift between ${\cal D}_{S}$ and ${\cal D}_{n}$ such that
${g_{n}}=\arg\mathop{\min}\limits_{{g_{n}}}\frac{1}{{{|{\cal D}_{n}|}}}\sum\limits_{j\in{{\cal D}_{n}}}{{\cal L}\left({{\bf{y}}_{j}^{n},{t_{S}}\left({{\bf{x}}_{j}^{n}}\right)}\right)}+\lambda{\ell_{{{\cal D}_{S}},{{\cal D}_{n}}}},$ (2)
where $\lambda>0$ controls the trade‑off, and $\ell_{\mathcal{D}_{S},\mathcal{D}_{n}}$ quantifies the discrepancy between domains, enabling personalized learning while keeping all sensitive data on‑device.
@@
Recent studies have explored TinyML–LargeML integration via TL across diverse applications. For example, TinyTL [19] uses ProxylessNAS-Mobile for source-domain pre-training and fine-tunes TinyML devices with a once-for-all network, applying parameter-based TL for tasks like object and facial attribute recognition of cars, aircraft, flowers, birds, pets, food, and celebrities. Similarly, [125] introduces a bidirectional integration using CNN pre-training with TensorFlow Lite Micro, pruning, and quantization for in-car presence detection and classification. In [131], a bidirectional integration system with feature-based TL leverages a deep neural network (DNN) source-domain backbone and a CMSIS-NN target model for IoT health trackers, supporting object and activity recognition. Furthermore, [12] applies feature-based TL with VGG-16, DenseNet, and MobileNet backbones, deployed through Edge Impulse’s EON Tuner for face mask detection. Another study [57] combines a DNN-long short-term memory (LSTM) model with parameter-based TL for soil humidity prediction in smart agriculture, fine-tuned on TinyML devices with TensorFlow Lite Micro and quantization. In [58], the authors introduce a feature-based TL bidirectional system for motion classification, using principal component analysis (PCA) for dimensionality reduction and a CNN-LSTM trained on large datasets, with Raspberry Pi as an edge intermediary platform. Finally, [84] presents a feature-based TL framework enhanced with meta-learning, using pre-trained MCUNet, MobileNetV2, and ProxylessNAS, then fine-tuned on TinyML devices via a task-adaptive sparse-update method to support few-shot cross-domain classification under dynamic constraints.
@@
IV-B Federated Transfer Learning (FTL)
TL methods (Section IV-A) work only when source and target domains share related tasks and some dataset similarities. In 6G systems, these conditions rarely hold because massive device populations generate highly non‑IID, heterogeneous, and context‑dependent data. FTL overcomes this by combining FL’s collaborative training with TL’s cross‑domain adaptation: source‑domain clients train the initial global model, and target‑domain clients then adapt and personalize it to their own data and tasks (see [197] for more details).
@@ IV-B1 Vanilla Federated Transfer Learning
Vanilla FTL is the core form of FTL in which multiple parties jointly train a model without sharing raw data, even when their feature spaces and data distributions differ. It links local feature extractors to a shared prediction model optimized through privacy‑preserving gradient exchange. This separation supports scalable knowledge reuse and fine‑grained personalization across heterogeneous 6G devices, making it well suited for privacy‑sensitive TinyML–LargeML deployments.
Fig. 5 illustrates the workflow of this FTL-based approach. The FTL process begins with the standard FL protocol, differing primarily in the initialization of the global model. Specifically, the FL phase includes the following steps.
@@
    Initial global model sharing: Unlike conventional FL, where the global model starts with random weights, FTL initializes it with a pre-trained LargeML model $f_{S}$, trained on the server’s large dataset ${\cal D}_{S}$ such that $f_{S}^{\left(0\right)}=\arg\mathop{\min}\limits_{{f_{S}}}\frac{1}{{{|{\cal D}_{S}|}}}\sum\nolimits_{i\in{{\cal D}_{S}}}{{\cal L}\left({{{\bf{y}}_{i}},{f_{S}}\left({{{\bf{x}}_{i}}}\right)}\right)}$. Once trained, $f_{S}^{\left(0\right)}$ is distributed to a source client set ${\cal K}=\left\{{1,\ldots,K}\right\}$.
@@
    Local training: In round $t$, each source client $k\in{\cal K}$} exploits its dataset ${\cal D}_{k}$} to train a local copy of the model using the loss function $f_{S,k}^{\left(t\right)}=\arg\mathop{\min}\nolimits_{f_{S}^{\left({t-1}\right)}}\frac{1}{|{\cal D}_{k}|}\sum\nolimits_{j\in{{\cal D}_{k}}}$ ${\cal L}\big({{\bf{y}}_{j}^{k},f_{S}^{\left({t-1}\right)}\left({{\bf{x}}_{j}^{k}}\right)}\big)$, adapting the model to local patterns.
@@
    Local model update: Subsequently, each client uploads its updated parameters $f_{S,k}^{\left(t\right)}$ to the server for aggregation.
@@
    Server-side aggregation: The server aggregates client updates into a new global model using either FedAvg [113], proportion‑based weighting [208], or advanced heterogeneity‑aware aggregation methods to minimize the global loss: $f_{S}^{\left(t\right)}=\arg\mathop{\min}_{f_{S,k}^{\left(t\right)}}\sum\nolimits_{k\in{\cal K}}\frac{r_{k}}{|{\cal D}_{k}|}\sum\nolimits_{j\in{{\cal D}_{k}}}$ ${\cal L}\big({{\bf{y}}_{j}^{k},f_{S,k}^{\left(t\right)}\left({{\bf{x}}_{j}^{k}}\right)}\big)$}, where ${r_{k}}={|{\cal D}_{k}|}/\sum\nolimits_{k=1}^{K}{{|{\cal D}_{k}|}}$ is the ratio of client $k$’s data size to the total data size.
@@
Step 6 presents a knowledge transfer phase. Once the global model has been sufficiently trained and demonstrates strong generalization, the server transfers the useful components of the pre-trained global model $f_{S}$, denoted $t_{S}\subset f_{S}$, to the target TinyML devices (target clients). This transferred model encapsulates the knowledge learned from source clients.
Step 7 stands for model personalization and on-device adaptation. Each target TinyML device $n\in{\cal N}=\left\{{1,\ldots,N}\right\}$ fine-tunes $t_{S}$ with its smaller, domain-specific dataset ${\cal D}_{n}$. Depending on task similarity, adaptation may involve updating all layers $t_{S}$ or only a subset. TinyML devices employ lightweight models $g_{n}$, trained with the objective function (2). This enables efficient personalization without retraining from scratch, conserving computational resources and reducing latency.
@@
Recent studies have applied FTL to integrate TinyML and LargeML across diverse domains. For example, TinyFedTL [80], a parameter-based FTL framework, pre-trains MobileNetV2 and adapts Perf-MobileNet (part of the TinyML Perf Benchmark) for personalized model adaptation on IoT edge devices, improving the quality of experience (QoE) under resource constraints without compromising privacy. FedHealth [26], a parameter-based FTL framework, employs a CNN backbone and fine-tuning for wearable healthcare, supporting activity recognition and auxiliary Parkinson’s disease diagnosis with high accuracy and privacy preservation. ACGAN-FTL [54] integrates auxiliary classifier generative adversarial networks for Raspberry Pi-based TinyML devices, predicting the quality of pre-baked carbon anodes in industrial production. In [48], a feature-based FTL bidirectional integration system using Genann outperformed standalone FL, TL, and TensorFlow Lite fusion across classification and regression tasks, demonstrating effective device-specific adaptation on Arduino WiFi Rev2, MKR1010, ESP8266, and ESP32.
@@
From a hierarchical FTL perspective, IoTDefender [43] employs three edge servers running FL, aggregated at a cloud server, with CNN-based pre-training and parameter-based TL for intrusion detection in resource-constrained IoT networks. Upon this, HFTL [132] introduces a hierarchical FTL framework that incorporates additional fog servers. These fog servers aggregate models from associated edge servers before forwarding them to a cloud server for further aggregation. Using VGG-16, ResNet50V2, and MobileNet as pre-trained models, fine-tuned with an additional DNN, HFTL achieves secure and efficient fault classification in additive manufacturing, with parameter-based TL is the core approach.